# You Name It, I Run It: An LLM Agent to Execute Tests of Arbitrary Projects

### v1

@@ 3.3. Costs
ExecutionAgent relies on a large language model, which incurs costs computed in terms of token usage. Figure 9 shows the monetary costs due to the usage of the LLM. Based on current pricing, the average cost of processing a project is 0.16 USD. The costs differs significantly depending on whether ExecutionAgent succeeds in building and testing the project. For projects that fail to build or test, the costs are higher because ExecutionAgent tries to fix any issued until exhausting the budget, leading to an average cost of 0.34 dollars. In contrast, for projects where the approach succeeds, the costs are lower, with an average of only 0.10 dollars. Overall, we consider the current costs to be acceptable given the benefits of the approach, and expect costs to decrease over time as LLMs become more efficient and cheaper.
@@
3.4. Ablation
To understand the importance of the two phases of ExecutionAgent, we conduct an ablation study featuring two variants of our approach: (1) ExecutionAgent without the preparation phase: In this variant, the approach starts directly with the feedback loop, i.e., without information obtained via meta-prompting and web search. (2) ExecutionAgent without the feedback loop: This variant replaces the feedback loop with a single query that provides the information collected in the preparation phase and asks the LLM to generate a full script that installs the given project and runs its tests. Due to budget limitations, we run theses variants on ten projects subsampled randomly while ensuring to include two projects per main language.
The results of these two variants, as well as the results of the full approach, are presented in Table 3. We find that the full approach is the most effective, with a successful build rate of 8/10 and a successful test execution rate of 7/10. In contrast, omitting either of the two phases leads to a significant decrease in effectiveness, with only 1 successfully executed test suites achieved by the no-feedback variant. Because the variants perform overall less work, they also incur lower costs. However, the moderate cost per project, along with the clear benefits of the full approach, show the value of the ExecutionAgent.
Omitting the preparation phase led to the agent not having a concrete plan and also ignoring best practices. Without the feedback loop, the Agent is similar to the LLM-scripts baseline but with more information and context in the prompt.
Table 3. Ablation study with variants of ExecutionAgent.
ExecutionAgent No preparation phase No feedback phase
Built/installed 8 / 10 1 / 10 7 / 10
Executed tests 7 / 10 0 / 10 1 / 10
Results similar to ground truth 7 / 10 0 / 10 1 / 10
Cost per project (USD) 0.12 0.10 0.01
@@ 3.5. Analysis and Discussion
3.5.1. Tools Usage
To understand the behavior of ExecutionAgent, we start by a quantitative analysis of its tools usage. Overall, 74.8% of calls made by the approach invoke the tool terminal, while 16.0% and 8.2% invoke write_file and read_file, respectively. The “end of task” tool accounts for 1% of all tool invocations. This high-level analysis shows that ExecutionAgent uses all available tools, and that the terminal seems to be indispensible.
To further uncover the commands executed through the terminal, we count the frequency of all invoked commands. Figure 10 shows the results, where “other” means all commands beyond those shown explicitly. We find that commands for exploring directories and files, such as ls, cd, and pwd, are common, summing up to 33%. Another subset of popular commands (21%) is installation and build commands, such as apt install for system packages, pip for python and npm for JavaScript. Finally, the high prevalence of “other” shows that ExecutionAgent uses a wide range of commands, suggesting that hard-coding a fixed set of commands would be ineffective.
@@ 3.5.3. Complexity of Scripts Generated by ExecutionAgent
We measure the script size (Section 3.1.2) to assess the complexity of the scripts generated by ExecutionAgent. Across the 50 projects, the average script size is 16, with a minimum of 7 and a maximum of 31. That is, the final setup requires running 16 commands, on average. While this number may not seem particularly high, the scripts are consise because the agents creates them once it has found a successful sequence of commands.
@@ 4. Threats to Validity
Our results are subject to several threats to validity. First, ExecutionAgent typically runs tests in a single configuration (e.g., one language version, browser, or operating system), which may not capture variations across different environments, such as multiple browsers or OS versions. This limitation can be addressed by allowing users to modify the prompt to specify the desired configurations. Second, the approach designed around modern technologies (e.g., as asked for in the meta-prompt), which may limit the effectiveness for projects that require older dependencies. However, this limitation is mitigated by the technique’s iterative approach, allowing it to detect and adapt to legacy dependencies when necessary. Finally, the approach has been primarily tested on popular projects with relatively good documentation. Results could differ on lesser-known or poorly documented projects.
@@ 6. Conclusion
In this paper, we introduced ExecutionAgent, an automated technique designed to address the complexities of test suite execution across diverse software projects. By leveraging a large language model-based agent, ExecutionAgent autonomously installs, configures, and runs tests on arbitrary projects, producing tailored scripts to streamline future test executions. Our approach is modeled after the decision-making processes of a human developer, using meta-prompting and iterative refinement to adapt to project-specific dependencies and configurations. An evaluation on 50 open-source projects using 14 languages demonstrates that ExecutionAgent successfully executes most test suites and closely matches ground truth results, with clear improvements over existing methods. With reasonable computational and financial costs, ExecutionAgent holds strong potential as a useful technique for developers, autonomous coding systems, and researchers needing reliable test execution capabilities across varied software ecosystems.
@@ 7. Data Availability
ExecutionAgent is available at https://github.com/sola-st/ExecutionAgent.

### v2

@@ 3.3. Costs
ExecutionAgent relies on an LLM, which incurs costs computed in terms of token usage. Figure 9 shows the monetary costs due to the usage of the LLM. Based on current pricing, the average cost of processing a project is USD 0.16. The costs differs significantly depending on whether ExecutionAgent succeeds in building and testing the project. For projects that fail to build or test, the costs are higher because ExecutionAgent tries to fix any issues until exhausting the budget, leading to an average cost of USD 0.34. In contrast, for projects where the approach succeeds, the costs are lower, with an average of only USD 0.10. Overall, we consider the current costs to be acceptable given the benefits of the approach, and expect costs to decrease over time as LLMs become more efficient and cheaper.
@@
3.4. Ablation Study
To evaluate the contributions of different components of ExecutionAgent, we conduct an ablation study with multiple variants. The studied variants are as follows:
  • No preparation phase: The agent starts directly with the feedback phase, lacking the structured information from meta-prompting and retrieval.
  • No feedback phase: Instead of iteratively refining its approach, the agent generates a single script using the information from the preparation phase.
  • No guidelines: The agent does not use language-specific or general guidelines, reducing its ability to follow common installation and testing conventions.
  • No external retrieval: No external information from web search or CI/CD scripts.
Table 3 presents the results of these variants compared to the full approach. The ablated versions demonstrate significantly lower performance. Removing the preparation phase leads to a notable drop in success rates (12 successful builds and only 8 test executions), highlighting the importance of structured planning and context retrieval. The absence of feedback further reduces effectiveness, as it eliminates the iterative refinement process, leading to only 5 successful test executions. We learn that generated guidelines play a crucial role, as without them, the effectiveness drops by 50%. While retrieving external information (e.g., CI/CD configurations and via web search) are important, removing it only drops the success rate by 21% and increases costs by 13%. Overall, these findings emphasize the importance of both the preparation phase and the feedback phase in ensuring successful project setup and testing.
Table 3. Ablation study with four variants of ExecutionAgent.
Built — Executed tests — Deviation ¡ 10% — Project cost (USD)
ExecutionAgent 41 33 29 0.16
No preparation phase 12 8 7 0.19
No feedback phase 19 5 5 0.02
No guidelines 26 15 15 0.15
No CI/Web retrieval 31 24 23 0.18
@@ 3.5. Analysis and Discussion
3.5.1. Usage of Tools
To understand the behavior of ExecutionAgent, we start by a quantitative analysis of its tool usage. Overall, 74.8% of calls made by the approach invoke the tool terminal, while 16.0% and 8.2% invoke write_file and read_file, respectively. The “end of task” tool accounts for 1% of all tool invocations. This high-level analysis shows that ExecutionAgent uses all available tools, and that the terminal seems to be indispensible.
To further uncover the commands executed through the terminal, we count the frequency of all invoked commands. Figure 10 shows the results for the top-10 commands. We find that commands for exploring directories and files, such as ls, cd, and pwd, are common, summing up to 33%. Another subset of popular commands (21%) is installation and build commands, such as apt install for system packages, pip for python and npm for JavaScript. Finally, approximately 25% of executed commands fall outside the top 10, demonstrating that ExecutionAgent utilizes a diverse command set, making a hard-coded selection of commands ineffective.
@@ 3.5.3. Complexity of Scripts Generated by ExecutionAgent
We measure the script size (Section 3.1.2) to assess the complexity of the scripts generated by ExecutionAgent. Across the 33 successful projects, the average script size is 18, with a minimum of 12 and a maximum of 37. That is, the final setup requires running 16 commands, on average. While this number may not seem particularly high, the scripts are concise because the agent creates them once it has found a successful sequence of commands.
@@ 4. Threats to Validity
Our results are subject to several threats to validity. First, ExecutionAgent typically runs tests in a single configuration (e.g., one language version, browser, or operating system), which may not capture variations across different environments, such as multiple browsers or OS versions. This limitation can be addressed by allowing users to modify the prompt to specify the desired configurations. Second, the approach is designed around modern technologies (e.g., as asked for in the meta-prompt), which may limit the effectiveness for projects that require older dependencies. However, this limitation is mitigated by the technique’s iterative approach, allowing it to detect and adapt to legacy dependencies when necessary. Third, the approach has been primarily tested on popular projects with relatively good documentation. Results could differ on less well-known or poorly documented projects. In addition, a lot of projects in our dataset do not use submodules, which seem to challenge ExecutionAgent. Finally, our criteria for selecting projects, especially the availability of ground truth test execution results, may bias the dataset toward projects that use some form of CI/CD. Addressing this potential bias would require another way of establishing a ground truth, such as manually performing the task addressed by ExecutionAgent and comparing the results to those produced by the approach.
@@ 6. Conclusion
In this paper, we introduced ExecutionAgent, an automated technique designed to address the complexities of test suite execution across diverse software projects. By leveraging a large language model-based agent, ExecutionAgent autonomously builds an arbitrary project and runs its tests, producing tailored scripts to streamline future test executions. Our approach is modeled after the decision-making processes of a human developer, using meta-prompting and iterative refinement to adapt to project-specific dependencies and configurations. An evaluation on 50 open-source projects using 14 languages demonstrates that ExecutionAgent successfully executes most test suites and closely matches ground truth results, with clear improvements over existing methods. With reasonable computational and financial costs, ExecutionAgent holds strong potential as a useful technique for developers, autonomous coding systems, and researchers needing reliable test execution capabilities across varied software ecosystems.
@@ 7. Data Availability
8. Acknowledgments
This work was supported by the European Research Council (ERC, grant agreements 851895 and 101155832) and by the German Research Foundation within the DeMoCo and QPTest projects.